# Investigation of $\Lambda_{c}$ States and $(\bar{D}N)$ Molecules Production at EicC and EIC

### v1

@@ I INTRODUCTION
Searching for hadronic molecule states in charm baryons is an interesting topic in hadron physics. Since the BaBar Collaboration observed $\Lambda_{c}(2940)$ in the $D^{0}p$ invariant mass distribution [1], as a likely candidate for a hadronic molecular state, many groups have studied its nature with different pictures.  He et al. [2] proposed that $\Lambda_{c}(2940)$ is a $D^{*0}p$ molecular state with the quantum number $J^{P}=1/2^{-}$. Using SU(4) effective Lagrangians,  Dong et al. [3] suggested $1/2^{+}$ for $\Lambda_{c}(2940)$ and excluded $1/2^{-}$ since the calculated partial widths are larger than the experimental results. In quark model calculations [4, 5, 6], the states closest in mass to $\Lambda_{c}(2940)$ are $\Lambda_{c}(\frac{1}{2}^{-},2P)$ and $\Lambda_{c}(\frac{3}{2}^{-},2P)$. Their masses are approximately 40 MeV and 60 MeV higher than that of $\Lambda_{c}(2940)$, respectively. In 2020,  Luo et al. [7] considered the coupled-channel effect by introducing the $D^{*}N$ contribution, and found that the mass of $\Lambda_{c}(\frac{3}{2}^{-},2P)$ could be consistent with $\Lambda_{c}(2940)$. However, in this scenario, the mass of $\Lambda_{c}(\frac{1}{2}^{-},2P)$ is higher than that of $\Lambda_{c}(\frac{3}{2}^{-},2P)$, which has never been observed before.
@@
Additionally, we also estimate the production rates of the predicted ($\bar{D}N$) molecular states [26] at EicC and EIC to study the feasibility of discovering these molecular states.
@@ II.3.1 Hadronic molecular structure
The Lagrangian of $(\bar{D}N)-\bar{D}N$ consists of two parts and we use $B^{\prime}$ and $B$ to represent $(\bar{D}N)$ and $N$.
@@
$\gamma_{5}\Gamma^{\mu}(q^{2})=\gamma_{5}[F_{1}(q^{2})\gamma^{\mu}+F_{2}(q^{2})p^{\mu}+F_{3}(q^{3})p^{\prime\mu}]$ (26)
@@ II.3.2 ${}^{3}P_{0}$ model
where $\alpha_{\rho}=0.4$ GeV. $m_{Q}$ and $m_{q}$ are the masses of the heavy and light quarks, respectively, with $m_{u}=m_{d}=220$ MeV, $m_{s}=419$ MeV, and $m_{c}=1628$ MeV. The harmonic oscillator parameters in mesons are $R=2.5\,\mathrm{GeV}^{-1}$ for light mesons and $R=1.67\,\mathrm{GeV}^{-1}$ for $D$ mesons [46, 47].
@@ III.2 $(\bar{D}N)$ states
Table 3: Coupling constants in $(\bar{D}N)-\bar{D}N$ vertex. The cutoff is set to $\Lambda=1$ GeV and $\Lambda_{1}=1$ GeV.

### v2

@@ I INTRODUCTION
Searching for hadronic molecule states in charm baryons is an interesting topic in hadron physics. Since the BaBar Collaboration observed $\Lambda_{c}(2940)$ in the $D^{0}p$ invariant mass distribution [1], as a likely candidate for a hadronic molecular state, many groups have studied its nature with different pictures.  He et al. [2] proposed that $\Lambda_{c}(2940)$ is a $D^{*0}p$ molecular state with the quantum number $J^{P}=1/2^{-}$. Using SU(4) effective Lagrangians,  Dong et al. [3] suggested $1/2^{+}$ for $\Lambda_{c}(2940)$ and excluded $1/2^{-}$ since the calculated partial widths are larger than the experimental results. In quark model calculations [4, 5, 6], the states closest in mass to $\Lambda_{c}(2940)$ are $\Lambda_{c}(\frac{1}{2}^{-},2P)$ and $\Lambda_{c}(\frac{3}{2}^{-},2P)$. Their masses are approximately 40 MeV and 60 MeV higher than that of $\Lambda_{c}(2940)$, respectively. In 2020,  Luo et al. [7] considered the coupled-channel effect by introducing the $D^{*}N$ contribution, and found that the mass of $\Lambda_{c}(\frac{3}{2}^{-},2P)$ could be consistent with $\Lambda_{c}(2940)$. However, in this scenario, the mass of $\Lambda_{c}(\frac{1}{2}^{-},2P)$ is higher than that of $\Lambda_{c}(\frac{3}{2}^{-},2P)$, which has never been observed before. While in Ref. [8], role of the $\Lambda_{c}(2940)$ in $\pi^{-}p\to D^{-}D^{0}p$ reaction was theoretical studied within an effective Lagrangian approach.
@@
Additionally, we also estimate the production rates of the predicted ($\bar{D}N$) molecular states [27, 28, 29] at EicC and EIC to study the feasibility of discovering these molecular states.
@@ II.3.1 Hadronic molecular structure
The Lagrangian of $(\bar{D}N)-\bar{D}^{*}N$ consists of two parts and we use $B^{\prime}$ and $B$ to represent $(\bar{D}N)$ and $N$.
@@
$\gamma_{5}\Gamma^{\mu}(q^{2})=\gamma_{5}[F_{1}(q^{2})\gamma^{\mu}+F_{2}(q^{2})p^{\mu}+F_{3}(q^{2})p^{\prime\mu}]$ (26)
@@ II.3.2 ${}^{3}P_{0}$ model
where $\alpha_{\rho}=0.4$ GeV. $m_{Q}$ and $m_{q}$ are the masses of the heavy and light quarks, respectively, with $m_{u}=m_{d}=220$ MeV, $m_{s}=419$ MeV, and $m_{c}=1628$ MeV. The harmonic oscillator parameters in mesons are $R=2.5\,\mathrm{GeV}^{-1}$ for light mesons, $R=1.67\,\mathrm{GeV}^{-1}$ for $D$ mesons, and $R=1.94\,\mathrm{GeV}^{-1}$ for $D^{*}$ mesons [49, 50, 47].
@@ III.2 $(\bar{D}N)$ states
Table 3: Coupling constants in $(\bar{D}N)-\bar{D}^{*}N$ vertex. The cutoff is set to $\Lambda=1$ GeV and $\Lambda_{1}=1$ GeV.